The sum of two numbers is 114. The second is 6 more than 5 times the first. What are the two numbers?
Answer:
One number is
And the other number is

Answer by Earlsdon(6294)

Let the two numbers be a and b.
a+b = 114 The sum of the two numbers is 114.
b = 5a+6 The second number is 6 more than 5 times the first number. Substitute this for b in the first equation and solve for a.
a + (5a+6) = 114 Simplify.
6a+6 = 114 Subtract 6 from both sides.
6a = 108 Divide both sides by 6.
a = 18
b = 5(18)+6
b = 96
One number is 18
The other number is 96